# Iterative Deepening Sampling for Large Language Models

### v1

@@ Abstract
The recent release of OpenAI’s o1 models and other similar frameworks showcasing test-time scaling laws has demonstrated their exceptional capability to tackle complex reasoning tasks. Inspired by this, subsequent research has revealed that such test-time scaling laws hinge on the model’s ability to search both within a single response (intra-response) and across multiple responses (inter-response) during training. Crucially, beyond selecting a single optimal response, the model must also develop robust self-correction capabilities within its own outputs. However, training models to achieve effective self-evaluation and self-correction remains a significant challenge, heavily dependent on the quality of self-reflection data. In this paper, we address this challenge by focusing on enhancing the quality of self-reflection data generation for complex problem-solving, which can subsequently improve the training of next-generation large language models (LLMs). Specifically, we explore how manually triggering a model’s self-correction mechanisms can improve performance on challenging reasoning tasks. To this end, we propose a novel iterative deepening sampling algorithm framework designed to enhance self-correction and generate higher-quality samples. Through extensive experiments on Math500 and AIME benchmarks, we demonstrate that our method achieves a higher success rate on difficult tasks and provide detailed ablation studies to analyze its effectiveness across diverse settings.
@@ 1 Introduction
Since ChatGPT, large language models (LLMs) have been a rapidly evolving domain that tries to solve problems beyond traditional language tasks like summarization or question answering (Chen et al., 2023; Yao et al., 2023; Chen et al., 2024b, d). Significantly, the newly released OpenAI O1 has demonstrated its strong capability in complex problem-solving through its detailed reasoning steps before outputting the final answer (Jaech et al., 2024). Since then, many researchers have studied how to replicate success from an open-source perspective and how to train models that are even better at efficiently solving problems that still remain unsolvable by the current LLMs (Huang et al., 2024; Zeng et al., 2024; DeepSeek-AI et al., 2025).
The remarkable capabilities of large language models (LLMs) have largely been driven by the pretraining scaling laws, which demonstrate that increasing the amount of data and model size leads to predictable improvements in performance. However, the availability of high-quality training data is inherently constrained—there is only one Internet from which to source such data. As a result, a growing research focus is on synthesizing high-quality data using existing models to further extend the limits of the pretraining scaling laws. Concurrently, following the release of OpenAI’s o1 series (Jaech et al., 2024), researchers have been exploring a new class of scaling laws that govern inference-time performance. These laws aim to optimize LLM performance given a larger computational budget at inference time, enabling improvements in complex problem-solving capabilities. Moreover, this paradigm not only enhances inference-time efficiency but also facilitates the generation of high-quality synthetic data, which can be leveraged to train the next generation of models and evaluate their performance (Guan et al., 2025).
@@
A key insight in this line of research is that high-quality data capable of eliciting self-reflection in LLMs is crucial for enhancing their reasoning capabilities. Notably, o1-like models demonstrate improved performance when longer reasoning chains are utilized (Huang et al., 2024; DeepSeek-AI et al., 2025). One straightforward yet effective approach to collecting such data involves manually inserting prompts such as ”Wait! Maybe I made some mistakes! I need to rethink from scratch.” These interventions have been shown to successfully trigger self-reflection, thereby improving the model’s reasoning depth and accuracy. However, as this remains an emerging area, many design choices in implementing self-reflection mechanisms lack systematic analysis. In particular, fundamental questions—such as when and how self-reflection should be introduced—remain largely unexplored.
@@
In this paper, we address the challenge of efficiently generating high-quality reasoning data using a fixed model, without additional training on complex reasoning tasks such as mathematical problem-solving. Specifically, we focus on manually introducing self-reflection triggers into generated data to improve the pass rate of a fixed model. To achieve this, we propose Iterative Deepening Sampling (ID-Sampling), a novel algorithmic framework that iteratively increases the sampling budget following a geometric progression, while incorporating self-reflection mechanisms at each expansion step. We theoretically demonstrate that ID-Sampling effectively balances computational efficiency and response quality, ensuring that the budget is not excessively wasted while still improving model performance. We evaluate ID-Sampling on two challenging reasoning benchmarks, MATH-500 and AIME-24, and demonstrate its effectiveness in Best-of-N sampling and majority voting settings. Additionally, we provide an ablation study analyzing how the rate of budget increase per iteration impacts both pass rate and inference time. Our results highlight the potential of ID-Sampling as a scalable approach for improving LLM reasoning performance through adaptive self-reflection mechanisms.
@@ 2 Related Works
Researchers have been exploring two primary directions to efficiently improve response quality under a frozen pre-trained LLM. One of the most commonly known technique is to increase the number of reasoning steps to encourage the LLMs to think more before the final output is provided (Wei et al., 2022). Following this direction, people have been using search to improve the performance of the inference of LLMs, and the recently released o1 models are one of the products that lead on this (Jaech et al., 2024). Following the categorization in Zeng et al. (2024), we define search as the process of LLMs making multiple attempts to solve the problem based on internal or external guidance, and the search strategy can be mainly categorized into tree-search and sequential revisions.
From the perspective of tree search, researchers try to generate a few responses at the same time and use external guidance to choose the one that is the best, i.e., tree search is searching inter-responses. Based on this classification, the simplest method in this direction is the Best-of-N (BoN) sampling (Cobbe et al., 2021; Brown et al., 2024). However, because of its efficiency compared to other actual search algorithms that involve a value estimation, there are many work that focuses on this setting to make it more efficient and effective. For example, Sun et al. (2024) proposed to use speculative rejection in BoN to reject bad responses through early score. Chen et al. (2024e) proposed to use extremely high temperature on the first token to greatly improve the Best-of-N performance on math and coding tasks. Besides BoN sampling, actual search that happens on the tree is also very popular. Inspired by the success of AlphaGo, many researchers believe Monte Carlo Tree Search can be a very efficient search algorithm to be combined with machine learning, and it is a recently hot topic in test-time computation as it has a more accurate value estimation and naturally balances exploration and exploitations (Zeng et al., 2024). Research for MCTS can be further classified into three classes based on how a node is defined, namely token-level actions (Zhang et al., 2023; Liu et al., 2023), step-level actions (Hao et al., 2023; Chen et al., 2024a; Zhou et al., 2023), and solution-level actions (Zhang et al., 2024a, b). On the other hand, beam search, as a very classic sampling algorithm, has been shown to be also efficient, and even better than MCTS in test-time scaling (Chen et al., 2024b). Similar to the different design in MCTS, Xie et al. (2024) proposed to let the LLM themselves be the value function and conduct beam search on a step level. TreeBoN(Qiu et al., 2024) proposed to use a similar step-level beam search data in a DPO training. Snell et al. (2024) proposed a k-step rollout to estimate the partial sequence in beam-search. In this paper, we adopt the simplest tree-search setting, Best-of-N (BoN) sampling, due to the absence of a highly accurate process reward model (PRM).
@@
The sequential revisions in the search strategy are more about self-evaluation and self-correction that happen sequentially in a single response, i.e, searching inner-responses. This direction is previously known as self-correction, and have also been known as iterations based on reflections. Starting from Self-Refine (Madaan et al., 2023), there are many works that studied how to effectively use LLM to give themselves their own evaluation and reflection (Yao et al., 2023; Shinn et al., 2023), and study how this reflection-based mechanism can be adapted to different applications (Chen et al., 2023; Gou et al., 2023; Chen et al., 2024c). While there have previously been a lot of discussions on whether LLM can actually self-evaluate and self-correct themselves (Huang et al., 2023; Chen et al., 2024f; Verma et al., 2024), recent studies have shown that by training with high-quality data that involve such process, LLM can actually achieve a useable level of capability and help self-correct in the process (Huang et al., 2024; Zeng et al., 2024; DeepSeek-AI et al., 2025).
While the research in the two search strategies is relatively separate, their methods are orthogonal and can be used together to make the sampling process more efficient (Snell et al., 2024), and there have also already been some trial that define a complete solution as a node in the tree search (Zhang et al., 2024a, b). This is especially important when generating training data for newer generations of models. In this paper, we focus on the intersection of both search strategies and study how to strategically trigger the sequential revision capability of LLMs more efficiently when tree search strategies are also used in the process.
@@ 3.1 Best-of-N Sampling
The BoN sampling is a simple yet effective method that can be fully parallelized to enhance performance. Increasing N guarantees improved results during training when a ground-truth checker is available and generally leads to better performance at inference time, provided that the reward model is sufficiently accurate. However, a key limitation of BoN is that all N samples are generated using the same sampling parameters, which restricts diversity in the output space. This lack of diversity often leads to performance saturation, preventing BoN from effectively scaling as N increases. Addressing this limitation remains an important research challenge for improving the robustness and scalability of BoN-based approaches.
@@ 3.2 Majority Voting
Similar to BoN sampling, majority voting, also known as self-consistency, provides an alternative approach for aggregating N different responses (Wang et al., 2023). As the name suggests, this method involves generating N responses, counting the frequency of each unique answer, and selecting the most frequently occurring response as the final output. This approach leverages the inherent redundancy in multiple generations to improve robustness and reliability, making it particularly useful for tasks requiring high confidence in correctness.
Vanilla majority voting offers the advantage of aggregating responses efficiently without relying on a reward model. Additionally, it can be extended to a weighted version, where weights are assigned based on PRM scores or confidence estimations of the generated answers (Wang et al., 2023). While majority voting benefits from not requiring a highly accurate reward model, it faces challenges in identifying equivalent answers in complex reasoning tasks. For instance, in mathematical problems, expressions such as $\frac{1}{\sqrt{3}}$ and $\frac{\sqrt{3}}{3}$ are equivalent but must be recognized as such to ensure correct vote counting. A common solution in mathematical domains involves using symbolic-based checkers to compare answer pairs and identify equivalences. However, this process can be computationally expensive, particularly when N responses exhibit high diversity, requiring up to $O(N^{2})$ comparisons. In some cases, this comparison overhead can be as time-consuming as the GPU inference itself, posing a significant bottleneck in large-scale reasoning tasks.
@@ 4 Iterative Deepening Sampling
1:  function ID-Sample(LLM, Dataset, $B_{0}$, $\gamma$):
@@
4:  while $Prefixes$ is not empty do
@@
6:     $Outputs\leftarrow\text{LLM.Search}(Prefixes,Budget)$
@@
10:        else
11:           $NewPrefix\leftarrow\textsc{PadTrigger}(Output)$
12:           $NewPrefixes.\text{append}(NewPrefix)$
13:        end if
14:     end for
15:     $Budget\leftarrow Budget\times\gamma$
16:     $Prefixes\leftarrow NewPrefixes$
17:  end while
18:  return $Prompt$
19:  end function
@@
In this paper, we explore a setting where tree-search-based strategies are combined with sequential revision. Given a fixed computational budget, an important challenge is determining how much additional budget should be allocated to refining a given prefix $x_{0}$ that the model has already sampled. Efficient budget allocation is crucial, as any saved resources can be redirected to increasing the number of N in Best-of-N sampling or deepening tree search, ultimately improving overall performance. Understanding this trade-off is key to optimizing both search efficiency and model output quality.
@@
More specifically, in the setting where sequential revision is considered, the self-evaluation and self-correction process can be manually triggered by introducing a predefined trigger sentence, such as ”Wait! Maybe I made some mistakes! I need to rethink from scratch.” which is used in our experiment. In most cases, LLMs respond to this trigger by restarting their reasoning process and making self-corrections. Repeatedly inserting this sentence increases the overall length of the reasoning trajectory, potentially improving problem-solving accuracy by facilitating iterative refinement. In this paper, we propose a method for strategically placing the trigger sentence at increasing intervals within longer contextual windows, aiming to balance computational efficiency and the effectiveness of self-correction.
@@
Suppose we have already used a budget of $b$ to generate a prefix $x_{0}$ and are now considering whether to immediately introduce a trigger sentence. Iterative Deepening (ID) sampling allocates an additional budget of $\gamma\dots b$ before inserting the next trigger sentence, where $\gamma>1$ is a tunable hyperparameter. This iterative process continues until reaching a maximum budget $B$, beyond which no further trigger sentences are introduced. If the reasoning process reaches a natural stopping point—i.e., a complete answer is generated within the allocated budget—the process terminates early. This is because generating a full response from scratch generally leads to more reliable outputs than attempting to refine an already complete solution. Additionally, continuing the reasoning process after an answer has been formed often requires the use of process reward models (PRMs), which tend to be less accurate in evaluating partial completions. Moreover, expecting an LLM to generate a different answer solely through self-correction assumes a highly robust self-revision capability, whereas the same computational cost could instead be used to generate an entirely new response from scratch. The complete procedure is outlined in Algorithm 1, where $B_{0}$ represents the initial budget. The function LLM.search performs tree search within a given budget and is adaptable to different tree-search strategies. The function PadTrigger handles the insertion of trigger sentences while ensuring redundancy is minimized if necessary.
The definition and allocation of computational budget depend on the specific tree-search framework employed, leading to variations in implementation strategies. In this paper, we focus on the setting where N responses are sampled in parallel, which we will discuss in detail later. We have provided an illustration of ID-sampling in Fig. 1.

### v2

# Iterative Deepening Sampling as Efficient Test-Time Scaling

@@ Abstract
Recent reasoning models, such as OpenAI’s O1 series, have demonstrated exceptional performance on complex reasoning tasks and revealed new test-time scaling laws. Inspired by this, many people have been studying how to train models to achieve effective self-evaluation and self-correction to further enable the scaling paradigm. However, less studied is how to efficiently scale test-time compute from a fixed model, and this remains a challenge. In this paper, we address this challenge by focusing on enhancing the quality of self-reflection data generation for complex problem-solving at test time, which can also subsequently improve the training of next-generation large language models (LLMs). Specifically, we explore how systematically triggering a model’s self-correction mechanisms can improve performance on challenging reasoning tasks. To this end, we propose a novel iterative deepening sampling algorithm framework designed to enhance self-correction and generate higher-quality samples. Through extensive experiments on Math500 and AIME benchmarks, we demonstrate that our method achieves a higher success rate on difficult tasks and provide detailed ablation studies to analyze its effectiveness across diverse settings.
@@ 1 Introduction
Since ChatGPT, large language models (LLMs) have been a rapidly evolving domain that tries to solve problems beyond traditional language tasks like summarization or question answering (Chen et al., 2023; Yao et al., 2023; Chen et al., 2024b, d). Significantly, the newly released OpenAI O1 has introduced its new paradigm of test-time scaling, which shows strong capability in complex problem-solving through its detailed reasoning steps before outputting the final answer (Jaech et al., 2024). Since then, many researchers have studied how to replicate success from an open-source perspective and how to train models that are even better at efficiently solving problems that still remain unsolvable by the current LLMs (Huang et al., 2024; Zeng et al., 2024; DeepSeek-AI et al., 2025). One key finding is that through reinforcement learning itself, LLM can spontaneously learn to self-evaluate and self-correct from time to time. However, there is no clear conclusion on whether self-evaluation is triggered often enough.
On the other hand, while training is the primary focus recently, it remains uncertain whether one could more efficiently scale test-time compute from a fixed model without additional changes in training or fine-tuning. Moreover, a more efficient sampling algorithm used at test-time can not only enhance inference-time efficiency but also facilitate the generation of high-quality synthetic data, which can be leveraged to train the next generation of models and evaluate their performance (Guan et al., 2025).
@@
A key insight in this line of research is that high-quality data capable of eliciting self-reflection in LLMs is crucial for enhancing their reasoning capabilities. Notably, o1-like models demonstrate improved performance when longer reasoning chains are utilized (Huang et al., 2024; DeepSeek-AI et al., 2025). One straightforward yet effective approach to collecting such data involves manually inserting prompts such as ”Wait! Maybe I made some mistakes! I need to rethink from scratch.” These interventions have been shown to successfully trigger self-reflection, thereby improving the model’s reasoning depth and accuracy. However, as this remains an emerging area, many design choices in implementing self-reflection mechanisms lack systematic analysis. In particular, fundamental questions, such as when and how self-reflection should be introduced, remain largely unexplored.
@@
In this paper, we address the challenge of efficient test-time scaling using a fixed model, without additional training, on complex reasoning tasks such as mathematical problem-solving. Specifically, we focus on manually introducing self-reflection triggers during the generation process to improve the pass rate of a fixed model. To achieve this, we propose Iterative Deepening Sampling (ID-Sampling), a novel algorithmic framework that iteratively increases the sampling budget following a geometric progression, while incorporating self-reflection mechanisms at each expansion step. We theoretically demonstrate that ID-Sampling effectively balances computational efficiency and response quality, ensuring that the budget is not excessively wasted while still improving model performance. We evaluate ID-Sampling on two challenging reasoning benchmarks, MATH-500 and AIME-24, and demonstrate its effectiveness in Best-of-N sampling and majority voting settings. Additionally, we provide an ablation study analyzing how the rate of budget increase per iteration impacts both pass rate and inference time. Our results highlight the potential of ID-Sampling as a scalable approach for improving LLM reasoning performance through adaptive self-reflection mechanisms.
@@ 2 Related Works
There are two primary directions for test-time scaling, namely scaling on multiple responses, and scaling the reasoning steps in a single response.
To efficiently scale to multiple responses, researchers have proposed both aggregating independent samples and employing structured methods such as tree-search-based architectures. While there have been many works on tree-search strategies early on (Zhang et al., 2023; Liu et al., 2023; Hao et al., 2023; Chen et al., 2024a; Zhou et al., 2023; Zhang et al., 2024a, b), researchers are putting more attention on aggregating independent responses given that it is hard to create a good process reward model to accurately estimate the quality of a partial solution (DeepSeek-AI et al., 2025). Specifically in this direction, researchers mostly rely on two main classes of aggregating strategies, namely self-consistency (Wang et al., 2022) and Best-of-N (Snell et al., 2024). There are many works that focus on this setting to make it more efficient and effective. For example, Sun et al. (Sun et al., 2024) proposed to use speculative rejection in BoN to reject bad responses through early scores. Chen et al. (Chen et al., 2024e) proposed to use extremely high temperature on the first token to greatly improve the Best-of-N performance on math and coding tasks.
@@
To scale within a single response, researchers have first introduced an intermediate step by the so-called chain-of-thoughts(Wang & Zhou, 2024). Then, starting with Self-Refine (Madaan et al., 2023), there are many works that studied how to effectively use an LLM to give themselves their own evaluation and reflection (Yao et al., 2023; Shinn et al., 2023), and study how this reflection-based mechanism can be adapted to different applications (Chen et al., 2023; Gou et al., 2023; Chen et al., 2024c). While there have previously been a lot of discussions on whether LLM can actually self-evaluate and self-correct themselves (Huang et al., 2023; Chen et al., 2024f; Verma et al., 2024), recent studies have shown that by training with high-quality data that involve such a process, LLM can actually achieve a useable level of capability and help self-correct in the process (Huang et al., 2024; Zeng et al., 2024). Inspired by the recent success of Deepseek-R1 (DeepSeek-AI et al., 2025), people have realized that scaling within a single-response can be efficiently achieved through reinforcement learning, and many follow-up works have also studied how to learn such scaling capability more efficiently (Wang et al., 2025). In this paper, we manually inject trigger sentence in the middle of generation, which is similar to the recent work s1 (Muennighoff et al., 2025). However, our work focuses on injecting at an early stage in the thinking process, while they focus on injecting when the current response completely finishes as a method for budget forcing.
While the research in the two search strategies is mostly separate, their methods are orthogonal and can be used together to make the sampling process more efficient (Snell et al., 2024), and some recent research has also studied how to balance the two, and has shown that scaling on the number of responses can be more efficient than scaling within a single response (Wang et al., 2025; Marjanović et al., 2025). However, their approach is empirically summarized based on a specific group of models distilled from Deepseek-R1. In this paper, we focus on the intersection of both scaling strategies and study how to strategically trigger the self-correction capability of LLMs more efficiently, considering scaling the number of responses as an option. We take into account the additional cost of scaling within one response by calculating the equivalent $N$ when comparing the answers.
@@ 3.1 Best-of-N Sampling
The BoN sampling is a simple yet effective method that can be fully parallelized to enhance performance. Increasing N guarantees improved results during training when a ground-truth checker is available and generally leads to better performance at inference time, provided that the reward model is sufficiently accurate. However, if paired with a reward model that is not good enough, BoN might fail to scale efficiently and might even decrease its performance when more samples are included for aggregation. Therefore, getting a good reward model is necessary for good performance for BoN.
@@ 3.2 Majority Voting
Similar to BoN sampling, majority voting, also known as self-consistency (cons@n), provides an alternative approach for aggregating N different responses (Wang et al., 2022). As the name suggests, this method involves generating N responses, counting the frequency of each unique answer, and selecting the most frequently occurring response as the final output. This approach leverages the inherent redundancy in multiple generations to improve robustness and reliability, making it particularly useful for tasks requiring high confidence in correctness.
Majority voting offers the advantage of aggregating responses efficiently without relying on a reward model. Additionally, it can be extended to a weighted version, where weights are assigned based on PRM scores or confidence estimations of the generated answers (Wang et al., 2022). While majority voting benefits from not requiring a highly accurate reward model, it faces challenges in identifying equivalent answers in complex reasoning tasks. For instance, it can be very hard to identify two code samples to be the same, and even in mathematical problems, expressions such as $\frac{1}{\sqrt{3}}$ and $\frac{\sqrt{3}}{3}$ are equivalent but must be recognized as such to ensure correct vote counting. A common solution in mathematical domains involves using symbolic-based checkers to compare answer pairs and identify equivalences. However, this process can be computationally expensive, requiring up to $O(N^{2})$ comparisons. In some cases, this comparison overhead can be as time-consuming as the GPU inference itself, posing a significant bottleneck in large-scale reasoning tasks.
@@ 4 Iterative Deepening Sampling
4.1 Motivation: Frequencies of Linguistic Markers in Reasoning Models
Figure 2: The frequency of linguistic markers related to thinking appeared in Deepseek-R1 on AIME-2024 dataset.
Recent studies have highlighted that overthinking can negatively impact the performance of reasoning models (Wang et al., 2025; Marjanović et al., 2025). A key observation is that correct answers tend to be shorter on average and contain fewer linguistic markers associated with ”thinking” behaviors.
Here, we present a novel perspective by analyzing not only the frequency but also the positional distribution of thinking-related markers within model responses. A complete list of these markers is provided in the appendix. As illustrated in Fig. 2, correct answers tend to exhibit fewer think markers, with a clear concentration in the early stages of the response. Their frequency drops sharply after the first 1000 tokens. In contrast, incorrect answers often contain self-corrections distributed throughout the response, with a noticeable decline only near the end. This distinct distribution pattern suggests that marker positioning may serve as a useful signal for assessing answer correctness.
4.2 Methods
@@
1:  function Generate_with_ID-Sampling(LLM, Dataset, $B_{0}$, $\gamma$, $B$):
@@
4:  while $Prefixes$ is not empty and $Budget\leq B$ do
@@
6:     $Outputs\leftarrow\text{LLM.generate}(input=Prefixes,max\_tokens=Budget)$
@@
10:        else if $Output$ finished thinking then
11:           $Output\leftarrow\text{LLM.generate}(input=Output)$
12:           LogAnswer($Output$)
13:        else
14:           $NewPrefix\leftarrow\textsc{PadTriggerSentence}(Output)$
15:           $NewPrefixes.\text{append}(NewPrefix)$
16:        end if
17:     end for
18:     $Budget\leftarrow Budget\times\gamma$
19:     $Prefixes\leftarrow NewPrefixes$
20:  end while
21:  end function
@@
In this paper, we focus on efficiently scaling test-time computation by replacing the standard sampling algorithm with a more effective, custom-designed alternative. Given a fixed overall computational budget, an important challenge is determining how much additional budget should be allocated to refining a given prefix $x_{0}$ that the model has already sampled. Efficient budget allocation is crucial, as any saved resources can be redirected to increasing the number of N in Best-of-N sampling or deepening tree search, ultimately improving overall performance. Understanding this trade-off is key to optimizing both search efficiency and model output quality.
@@
More specifically, the self-evaluation and self-correction process of the LLM can be manually triggered by introducing a predefined trigger sentence, such as ”Wait! Maybe I made some mistakes! I need to rethink from scratch.” or simply ”Wait”. In most cases, LLMs respond to this trigger by restarting their reasoning process and making self-corrections. Repeatedly inserting this sentence increases the overall length of the reasoning trajectory, potentially improving problem-solving accuracy by facilitating iterative refinement. In this paper, we propose a method for strategically placing a fixed trigger sentence at increasing intervals within longer contextual windows, aiming to balance computational efficiency and the effectiveness of self-correction.
@@
Motivated by the analysis in Section 4.1, we propose to introduce more trigger sentences at earlier stages and gradually reduce their occurrence as the response length increases. Suppose we have already used a budget of $b$ to generate a prefix $x_{0}$ and are now considering whether to immediately introduce a trigger sentence. Iterative Deepening (ID) sampling allocates an additional budget of $\gamma\times b$ before inserting the trigger sentence the next time, where $\gamma>1$ is a tunable hyperparameter. This iterative process continues until reaching a maximum budget $B$, beyond which no further trigger sentences are introduced. If the reasoning process reaches a natural stopping point—i.e., a complete answer is generated within the allocated budget—the process terminates early. This is because generating a full response from scratch generally leads to more reliable outputs than attempting to refine an already complete solution. The complete procedure is outlined in Algorithm 1, where $B_{0}$ represents the initial budget. The function LLM.generate conducts the generation within a given budget and is adaptable to different tree-search strategies. The function PadTriggerSentence handles the insertion of trigger sentences while ensuring redundancy is minimized if necessary.
[…]
Leveraging modern serving engines to efficiently allocate the KV-cache and execute inference in batches using data parallelism and pipeline parallelism significantly enhances the efficiency of ID-sampling. By setting a maximum generation token limit in the sampling parameters, the computational cost still scales linearly with generation time, ensuring that ID-sampling remains efficient even as the number of iterations increases.
4.3 Theoretical Analysis